OpenStudy (anonymous):

Braden's family had 1 and one-third gallons of milk in the refrigerator before dinner. After dinner, they had finished three-eighths of the milk. How many gallons of milk did they drink at dinner?

OpenStudy (anonymous):

\[total milk=1+\frac{ 1 }{ 3}=\frac{ 3+1 }{3 }=\frac{ 4 }{ 3 }\] milk drank after dinner\[=\frac{ 4 }{3 }\times \frac{ 3 }{8 }=?\]

OpenStudy (anonymous):

12/24 = 1/2

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

yw.